# Group-in-Group Policy Optimization for LLM Agent Training

### v2

@@ 5.2 Performance on ALFWorld and WebShop
Table 1 demonstrates the strong performance of GiGPO across both ALFWorld and WebShop. As shown, closed-source LLMs offer only moderate performance: Gemini-2.5-Pro reaches 60.3% success on ALFWorld and 35.9% on WebShop, while GPT-4o lags further behind. Open-source prompt-only agents (e.g., ReAct and Reflexion) show marginal improvements over vanilla prompting but still underperform, underscoring the difficulty of long-horizon control without post-training. RL training brings substantial gains: PPO improves average ALFWorld success to 54.4% on the 1.5B model and 80.4% on the 7B model, with WebShop scores also increasing significantly. However, this comes at the expense of increased complexity: requiring a separate critic network, hyperparameter tuning, and longer training durations [71, 49]. GRPO and RLOO also yield strong performance while being more computationally efficient, demonstrating the effectiveness of group-based RL in large-scale LLM training. Nevertheless, their lack of fine-grained per-step feedback limits their ability to provide precise credit assignment across long horizons. In contrast, GiGPO overcomes this limitation with a two-level advantage estimation, enabling both GiGPOw/ std and GiGPOw/o std to consistently surpass GRPO and RLOO. In particular, GiGPOw/o std surpasses GRPO by 13.3% on ALFWorld and 10.6% on WebShop at 1.5B, and by 12.6% and 9.1%, respectively, at 7B. These results highlight GiGPO’s superior ability to train LLM agents more effectively and more efficiently. We also find that GiGPO enables agents to exhibit emergent reasoning behavior (see Appendix E).
Lastly, we observe that the normalization factor $F_{\text{norm}}$ is task-dependent rather than universally helpful. On relatively difficult tasks (such as Look, Pick2, and WebShop), standard-deviation scaling ($F_{\text{norm}}=\text{std}$) could exaggerate gradients from overly difficult samples or highly imbalanced groups, harming update stability; fixing $F_{\text{norm}}=1$ therefore yields higher success. Yet, $F_{\text{norm}}=1$ offers no clear advantage on other tasks and both variants perform similarly, which aligns with findings in [72]. This suggests that $F_{\text{norm}}=\text{std}$ can still be beneficial when reward variance is stable.
@@ 5.3 Performance on QA tasks
As shown in Table 2, GiGPO achieves strong and consistent gains on multi-turn search-augmented QA tasks, reaching 42.1% at 3B and 47.2% at 7B, and significantly outperforming prior strong baselines such as Search-R1 and StepSearch. Although search-augmented QA is relatively short-horizon, the step-level signals of GiGPO still yield meaningful improvements. Furthermore, we observe that GiGPO is markedly more tool-efficient. Under a limit of at most 3 tool calls per query, the 7B model requires only $\sim$0.9 calls on average for single-hop tasks and $\sim$1.6 calls on average for multi-hop tasks: matching the superior performance of OTC [61], which achieves $\sim$1.0 and $\sim$1.7 calls, respectively, on the same tasks. This efficiency likely stems from GiGPO’s ability to effectively identify and suppress redundant queries in multi-turn decision-making. For instance, in repetitive patterns such as query1 $\rightarrow$ info1 $\rightarrow$ query1 $\rightarrow$ info1 $\rightarrow$ query1 $\rightarrow$ info1, GiGPO groups them into the same step-level group and eliminates unnecessary repeated calls during the training, similar to the example we introduced in Figure 3.
@@ 5.5 Dynamics of Step-Level Group
As shown in Figure 5, we observe that step-level groups of size 1 (i.e., those with $|G^{S}(\tilde{\bm{s}})|=1$) only account for < 35% throughout training. This indicates that the majority of states (over 65%) recur across trajectories and therefore contribute to achor state grouping. Moreover, at iteration 10, large group sizes $|G^{S}(\tilde{\bm{s}})|\geq 10$ account for over 20%, reflecting behavioral redundancy in the early stages of training. This is consistent with the fact that immature policies often produce invalid actions or fall into repetitive loops. As training progresses, the group size distribution changes markedly. By iteration 75, we observe a substantial reduction in extreme group sizes: $10\leq|G^{S}(\tilde{\bm{s}})|<50$ drops from 16.2% to 12.1% and $|G^{S}(\tilde{\bm{s}})|\geq 50$ drops from 5.6% to 3.1%. This shift suggests that the agent is learning to avoid previously common dead ends and invalid actions, and is beginning to exhibit more diverse and purposeful decision-making. At iteration 140, the distribution becomes tightly concentrated around group sizes of 6 to 8. Given that $N$ is set to $8$, this convergence implies that all 8 trajectories within an episode group are now behaving consistently: LLM agent has learned a coherent and robust policy for completing the task, aligning with the plateau in success rate ($80\%$).
@@ 6 Conclusions and Limitations
Appendix A Source Code
As part of the new assets released with this work, we provide the complete source code at https://github.com/langfengQ/verl-agent. The repository includes a comprehensive README.md with installation instructions, training and evaluation procedures, and license information.
Our code is built upon the veRL framework [73] and extended with several features to support scalable training of long-horizon LLM agents. Key capabilities of our code include: (1) a step-wise multi-turn interaction paradigm that avoids concatenating the entire interaction history into a single input, as done in RAGEN [51], thereby ensuring scalability for very long-horizon optimization, (2) support for parallelized, gym-style group-based environments to enable high-throughput and state-aligned rollouts, (3) a rich suite of environments, including Search (tool use), ALFWorld, WebShop, Sokoban, and Gym Cards; and (4) compatibility with both text-based and vision-language agents.
We provide a detailed installation guide (setup takes approximately five minutes), along with out-of-the-box scripts to reproduce our main results in ALFWorld and WebShop, as well as additional experiments on Sokoban and Gym Cards presented in Appendix D.3.
Appendix B Unbiasedness
@@
Appendix C Pseudo Code
@@
Appendix D Experiment Details
D.1 Details of Training
@@ D.2 Prompts
Prompt Template for WebShop You are an expert autonomous agent operating in the WebShop e-commerce environment. Your task is to: {task_description}. Prior to this step, you have already taken {step_count} step(s). Below are the most recent {history_length} observations and the corresponding actions you took: {action_history}. You are now at step {current_step} and your current observation is: {current_observation}. Your admissible actions for the current situation are: [{available_actions}]. Now it’s your turn to take one action for the current step. You should first reason step-by-step about the current situation, then think carefully which admissible action best advances the shopping goal. This reasoning process MUST be enclosed within <think> </think> tags. Once you’ve finished your reasoning, you should choose an admissible action for current step and present it within <action> </action> tags.
@@
D.3 Performance on Vision-Language Agents
@@
D.4 Orthogonality to Single-Turn Group-Based RL
@@
Appendix E Reasoning Behavior

### v3

@@ 5.2 Performance on ALFWorld and WebShop
Table 1 demonstrates the strong performance of GiGPO across both ALFWorld and WebShop. As shown, closed-source LLMs offer only moderate performance: Gemini-2.5-Pro reaches 60.3% success on ALFWorld and 35.9% on WebShop, while GPT-4o lags further behind. Open-source prompt-only agents (e.g., ReAct and Reflexion) show marginal improvements over vanilla prompting but still underperform, underscoring the difficulty of long-horizon control without post-training. RL training brings substantial gains: PPO improves average ALFWorld success to 54.4% on the 1.5B model and 80.4% on the 7B model, with WebShop scores also increasing significantly. However, this comes at the expense of increased complexity: requiring a separate critic network, hyperparameter tuning, and longer training durations kazemnejad2024vineppo ; chen2025reinforcement . GRPO and RLOO also yield strong performance while being more computationally efficient, demonstrating the effectiveness of group-based RL in large-scale LLM training. Nevertheless, their lack of fine-grained per-step feedback limits their ability to provide precise credit assignment across long horizons. In contrast, GiGPO overcomes this limitation with a two-level advantage estimation, enabling both GiGPOw/ std and GiGPOw/o std to consistently surpass GRPO and RLOO. In particular, GiGPOw/o std surpasses GRPO by 13.3% on ALFWorld and 10.6% on WebShop at 1.5B, and by 12.6% and 9.1%, respectively, at 7B. These results highlight GiGPO’s superior ability to train LLM agents more effectively and more efficiently. We also find that GiGPO enables agents to exhibit emergent reasoning behavior (see Appendix F).
Lastly, we observe that the normalization factor $F_{\text{norm}}$ is task-dependent rather than universally helpful. On relatively difficult tasks (such as Look, Pick2, and WebShop), standard-deviation scaling ($F_{\text{norm}}=\text{std}$) could exaggerate gradients from overly difficult samples or highly imbalanced groups, harming update stability; fixing $F_{\text{norm}}=1$ therefore yields higher success. Yet, $F_{\text{norm}}=1$ offers no clear advantage on other tasks and both variants perform similarly, which aligns with findings in chu2025gpg . This suggests that $F_{\text{norm}}=\text{std}$ can still be beneficial when reward variance is stable.
@@ 5.3 Performance on QA tasks
As shown in Table 2, GiGPO achieves strong and consistent gains on multi-turn search-augmented QA tasks, reaching 42.1% at 3B and 47.2% at 7B, and significantly outperforming prior strong baselines such as Search-R1 and StepSearch. Although search-augmented QA is relatively short-horizon, the step-level signals of GiGPO still yield meaningful improvements. Furthermore, we observe that GiGPO is markedly more tool-efficient. Under a limit of at most 3 tool calls per query, the 7B model requires only $\sim$0.9 calls on average for single-hop tasks and $\sim$1.6 calls on average for multi-hop tasks: matching the superior performance of OTC wang2025otc , which achieves $\sim$1.0 and $\sim$1.7 calls, respectively, on the same tasks. This efficiency likely stems from GiGPO’s ability to effectively identify and suppress redundant queries in multi-turn decision-making. For instance, in repetitive patterns such as query1 $\rightarrow$ info1 $\rightarrow$ query1 $\rightarrow$ info1 $\rightarrow$ query1 $\rightarrow$ info1, GiGPO groups them into the same step-level group and eliminates unnecessary repeated calls during the training, similar to the example we introduced in Figure 3.
@@ 5.5 Dynamics of Step-Level Group
As shown in Figure 5, we observe that step-level groups of size 1 (i.e., those with $|G^{S}(\tilde{\bm{s}})|=1$) only account for < 35% throughout training. This indicates that the majority of states (over 65%) recur across trajectories and therefore contribute to anchor state grouping. Moreover, at iteration 10, large group sizes $|G^{S}(\tilde{\bm{s}})|\geq 10$ account for over 20%, reflecting behavioral redundancy in the early stages of training. This is consistent with the fact that immature policies often produce invalid actions or fall into repetitive loops. As training progresses, the group size distribution changes markedly. By iteration 75, we observe a substantial reduction in extreme group sizes: $10\leq|G^{S}(\tilde{\bm{s}})|<50$ drops from 16.2% to 12.1% and $|G^{S}(\tilde{\bm{s}})|\geq 50$ drops from 5.6% to 3.1%. This shift suggests that the agent is learning to avoid previously common dead ends and invalid actions, and is beginning to exhibit more diverse and purposeful decision-making. At iteration 140, the distribution becomes tightly concentrated around group sizes of 6 to 8. Given that $N$ is set to $8$, this convergence implies that all 8 trajectories within an episode group are now behaving consistently: LLM agent has learned a coherent and robust policy for completing the task, aligning with the plateau in success rate ($80\%$).
@@ 6 Conclusions and Limitations
Acknowledgements
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG18/24).
Appendix A Open Source Codebase: verl-agent
As part of the new assets released with this work, we propose verl-agent (https://github.com/langfengQ/verl-agent), a highly scalable RL training framework for long-horizon, multi-turn LLM agent training.
Our verl-agent is built upon the veRL framework [73] and extends it with several features to enable scalable reinforcement learning for long-horizon LLM agents. Key capabilities of our framework include: (1) step-wise multi-turn interaction paradigm that avoids concatenating full interaction histories (as in Search-R1 [58] and RAGEN [51]), ensuring efficient memory control and scalability for very long-horizon optimization; (2) customizable memory module that allows developers to flexibly determine which historical information to include at each step (e.g., key events, summaries, or external knowledge); (3) parallel and group-based environments with a gym-style interface supporting high-throughput rollouts; (4) broad model compatibility, including Qwen3 [74], Qwen2.5, and LLaMA3.2, along with LoRA-based fine-tuning [75] for efficient large-model adaptation; (5) support for multimodal (vision-language) agents such as Qwen2.5-VL; (6) a diverse suite of environments, including Search (tool use), ALFWorld, WebShop, Sokoban, and Gym Cards; (7) comprehensive RL algorithm support, encompassing GiGPO, GRPO, PPO, DAPO, RLOO, etc.
Figure 7: Open-source agentic training framework comparison. Left: RAGEN/Search-R1 concatenates the full history at every step, leading to rapidly expanding context. Right: verl-agent adopts a step-wise multi-turn rollout with flexible per-step input construction and memory control.
Appendix B Broader Impacts
This work introduces an efficient, group-based RL algorithm, GiGPO, that enables fine-grained credit assignment for long-horizon LLM agents. By combining episode-level and step-level advantages, GiGPO improves training stability and agent performance without increasing computational or memory overhead.
The algorithm holds promise for a wide range of applications, including virtual assistants, web automation, educational tools, and embodied AI systems. Its critic-free and scalable design lowers the barrier to training effective multi-step agents, making advanced agent capabilities more accessible to both academic and industrial practitioners. While GiGPO is a methodological contribution with no direct downstream deployment, the improved agent training techniques may indirectly enable more autonomous behavior. As with any advancement in agent capabilities, it is important to consider appropriate safeguards and responsible usage. Overall, GiGPO contributes a valuable tool to the RL and LLM communities, with the potential to accelerate progress in building more intelligent, efficient, and adaptable AI agents.
Appendix C Unbiasedness
@@
Appendix D Pseudo Code
@@
Appendix E Experiment Details
E.1 Details of Training
@@ E.2 Prompts
Prompt Template for WebShop You are an expert autonomous agent operating in the WebShop e‑commerce environment. Your task is to: {task_description}. Prior to this step, you have already taken {step_count} step(s). Below are the most recent {history_length} observations and the corresponding actions you took: {action_history}. You are now at step {current_step} and your current observation is: {current_observation}. Your admissible actions for the current situation are: [{available_actions}]. Now it’s your turn to take one action for the current step. You should first reason step-by-step about the current situation, then think carefully which admissible action best advances the shopping goal. This reasoning process MUST be enclosed within <think> </think> tags. Once you’ve finished your reasoning, you should choose an admissible action for current step and present it within <action> </action> tags.
@@
E.3 Performance on Vision-Language Agents
@@
E.4 Orthogonality to Single-Turn Group-Based RL
@@
E.5 Sensitivity Analysis on $\omega$
We further analyze GiGPO’s sensitivity to the key hyperparameter $\omega$, which balances the episode-level advantage $A^{E}$ and the step-level advantage $A^{S}$. We conducted experiments on the WebShop task, using Qwen2.5-1.5B-Instruct.
Table 5: Sensitivity analysis on $\omega$ for the WebShop task.
$\omega$ 0.0 0.2 0.4 0.6 0.8 1.0 1.2 1.4
Score 76.2 79.6 82.4 83.5 84.9 83.5 82.6 77.0
Success Rate (%) 56.6 63.1 65.8 67.2 68.3 67.4 66.5 56.3
As shown in Table 5, GiGPO needs an appropriate $\omega$ to work best. Increasing $\omega$ initially improves performance due to the added fine-grained step-level reward. However, performance declines beyond the optimum ($\omega=0.8$), suggesting that excessive emphasis on step-level signals can suppress useful trajectory-level guidance. Furthermore, GiGPO is relatively insensitive to $\omega$ within the range $[0.4,1.2]$, demonstrating a reasonable degree of robustness to $\omega$.
Appendix F Reasoning Behavior